# Optical Image-to-Image Translation Using Denoising Diffusion Models: Heterogeneous Change Detection as a Use Case

### v1

@@ II-B Inference
The inference procedure, illustrated by Figure 3, aims at maximizing the image generation quality by selecting the best possible initial noise matrix to be fed to the reverse diffusion process. We adapt the classifier-free guided DDIM approach by integrating the same color standardization methods used during training. To further enhance the fidelity of the generated patches to their inputs, we refine the selection process for the initial noise condition input for the reverse diffusion process, ensuring optimal consistency and translation quality.
In our experiments, we observed that, for the same conditioning input $\bm{x}$, the quality of DDIM reverse diffusion outputs generated with different initial noise matrices varied significantly. That is, for a combination of independent and identically distributed random initial noise matrices, fixing $\bm{x}$, the DDIM, in some cases, generates samples entirely consistent with $\bm{x}$, whereas sometimes it generates inconsistent samples with undesired hallucinations. Since the DDIM reverse diffusion is a deterministic process, i.e., given fixed $\bm{\varepsilon}_{T}$ and $\bm{x}$, its output remains the same, the cause for varying consistency levels lies only in the different sampled initial noise matrices, which are the source of the randomness of the whole inference process. Thus, we propose a procedure that seeks to select an initial noise matrix such that the DDIM output possesses the highest consistency compared to $\bm{x}$. This procedure consists of repeating the DDIM reverse diffusion process with a reduced number of iterations, using as input $n_{\text{noisy}}$ different initial noise matrices. The resulting outputs are tested for consistency by calculating the peak signal-to-noise ratio (PSNR) between $\bm{x}_{\text{local}}$ and each noise matrix individually. The one that was responsible for the maximum PSNR is used as the definitive starting noise matrix for the inference process, which takes place with a higher number of iterations, to ensure maximal generation quality.
@@
III Model Setup
TABLE I: Hyperparameter values and their purpose for the performed experiments.
Parameter Value Purpose
$N_{\text{DDIM}}$ 64 Number of DDIM iterations for the final reverse diffusion in Line 16 of Alg. 3.
$d$ 8 Constant used to reduce the number of DDIM iterations for the $\bm{\varepsilon}_{T}^{\text{best}}$ estimation in Line 15 of Alg. 3.
$T$ $1024$ Number of forward diffusion noising steps.
$\varepsilon_{\text{consist}}$ $10$ It tells how far from the sampled timestep $t$ can the consistency timestep $t^{\prime}$ be sampled [Daras2023]. That is, $t^{\prime}\in\{t-\varepsilon_{\text{consist}},\dots,t+\varepsilon_{\text{consist}}\}$.
The model chosen for the experiments is the one presented in [Wang2022], where the authors propose a U-Net-style model that uses attention mechanisms to improve image fidelity. Differently from the diffusion-based models presented in [Saharia2021, Saharia2021SR3, ho2022classifierfree], this model processes the noisy input $\tilde{\bm{y}}$ and the conditioning input $\bm{x}$ independently. That is, $\tilde{\bm{y}}$ is fed right into the encoder and is processed normally, whereas $\bm{x}$ is re-introduced to the decoder by inserting it into each residual block. To properly match the spatial resolution required by each residual block, $\bm{x}$ is processed by spatially adaptive normalization (SPADE) blocks. As stated by the authors of the model, we observed that this modification improved significantly the generation quality and fidelity to the conditioning input. Some internal parameters of the model were adjusted to maximize performance while limiting its size due to hardware constraints. The number of channels has been set to be a multiple of $96$. The number of residual blocks for each downsampling operation has been set to $2$. Attention resolutions were set to $2,4,8$ and $16$. Finally, the number of attention heads has been set to $4$. For training, batch size has been set to $5$, and the learning rate has been set to $10^{-4}$. We use the Huber loss function [meyer2020alternative] with its parameter $\delta_{\ell}$ set to $0.5$ since it has been shown to reduce the impact of outliers in the gradients while maintaining the benefits of the mean square distance for non-outliers. RAdam optimizer [liu2021variance] has been chosen due to its proven training speed and its low sensitivity to the learning rate choice. The training ran for $31$ epochs, which was more than enough for the loss function to converge completely.
To stabilize training and avoid catastrophic forgetting of features, Exponential Moving Average (EMA) has been used in several DDM-based methods [Ho2020, Saharia2021, Wang2022, Dhariwal2021]. With the same goal, we instead use Stochastic Weight Averaging (SWA) [Izmailov2018], which consists of averaging weights after the training loss shows convergence behavior. Differently from EMA, SWA does not give higher importance to the weights of more recent epochs, but instead considers all epochs after the beginning of weight averaging to have the same importance for the overall learning. Since we have observed that the loss function converged after a number of epochs, it was reasonable to consider that newer weight updates would not bring an overall improvement of the model, thus applying SWA instead of EMA made more sense. The work presented in [garipov2018loss] suggests that averaging weights of different epochs can indeed result in higher performance than when only using the last weights of a training experiment. We observed that training stabilized after 10 epochs, so we started the SWA process at this moment. Also, SWA suggests increasing the learning rate during the epochs where weight averaging takes place. However, we noticed that increasing the learning rate over $10^{-4}$ made training unstable. For this reason, it is kept constant during the whole process.
For the performed experiments, multiple hyperparameters not directly related to the model setup have been chosen. They are presented in Table I, along with their purpose. Some choices were based on a grid search, and others were based on the articles that defined them in the first place.
Two recently proposed training techniques for denoising diffusion models were included in our experiments: Consistency Diffusion Loss and Min-SNR-$\gamma$ Weighting.
Consistency Diffusion Loss
It is a method to mitigate sampling drift in diffusion models by learning to be consistent [Daras2023]. It is based on enforcing a consistency property, which states that predictions of the model must be consistent across time when the conditioning input is the same. We observed that this significantly improved the consistency between neighbor patches of the same image frame and that it helped to reduce the frequency of undesired model hallucinations. The parameters related to consistency loss are:
  • $\lambda_{\text{consist}}$ This is the weight of the consistency loss over the final loss. It controls how much the consistency loss is used for learning relative to the DDPM loss.
  • $n_{\text{consist}}$: This is the number of consistency training steps to be executed in a single overall training step.
Min-SNR-$\gamma$ Weighting
It is a method to improve the convergence speed and effectiveness of DDPM training by adapting the loss weights of different timesteps of the diffusion process according to their signal-to-noise ratios (SNRs) [Hang2023]. It is based on the observation that the diffusion training suffers from conflicting optimization directions between timesteps, which slows down the learning process, as also visualized in our experiments. The single parameter of this method, $\gamma$, which hereby is referred to as $\gamma_{\text{SNR}}$, controls the degree of clamping the SNRs, which maintains stability and speed of the training process.
IV Dataset
@@
For the proposed experiments, paired images from satellites Sentinel-II and Planet Dove were gathered. The former is a widely known open-source optical satellite with a spatial resolution of 10 meters, whereas the latter is a class of commercial satellites aimed at producing high-quality images at a spatial resolution of 3 meters. A dataset containing 77 pairs of rasters has been assembled for training. These are parts of or a combination of acquisitions that were captured at similar time periods. They are properly geolocated and checked for image quality issues. In total, a vast area of 7894 $7894\text{ km}^{2}$ is covered by the training set, which is comparable to the area of Puerto Rico (9104 $9104\text{ km}^{2}$). As the test set for the image-to-image translation algorithm, 8 pairs of images were selected, which, in total, cover an area of 707 $707\text{ km}^{2}$. With these, image quality metrics are extracted for the proposed method and compared to other possible approaches for the I2I translation task.
@@
For evaluation of the performance of the method in a change detection task, we use pre- and post-images from the port region in Beirut, Lebanon. On August 4th, 2020, an explosion dramatically affected the port and surrounding areas. As they have recovered over time, many changes took place, e.g., new and repaired constructions and structures. Thus, for the change detection test, we use as pre-event images a pair of images acquired at a time close to the explosion, between 04 and 20 of August 2020, when the effects of the event are still evident. As post images, acquisitions from July 2023 of the exact same region are used. Figure 1 shows the Beirut pre and post-event images from Sentinel-II and Planet Dove. We also perform change detection tests in a region in Austin, USA, where many urban changes can be observed between the selected pre- and post-event images. They were captured in July 2021 and 2023, respectively.
@@
(b) Regression
(c) Regression + $\mathbf{W},\mathbf{C}$
(d) DDIM
(e) DDIM + $\mathbf{W},\mathbf{C}$
(f) DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting
Figure 4: Comparison among images generated by the tested I2I methods (b)-(f) and the original Sentinel-II image (a). It is evident how the image generated with the proposed model (f) displays higher resolution features not observable among the others, while maintaining patch-wise feature consistency.
V Experiments
V-A Image Generation
V-A1 Image Quality Metrics and Ablation Experiments
TABLE II: Image Quality Metrics of the Proposed Algorithm and Compared Approaches
@@ V-A1 Image Quality Metrics and Ablation Experiments
Sentinel-II 0.2977 98.08 12.76
Regression 0.2797 64.12 16.72
Regression + $\mathbf{W},\mathbf{C}$ 0.2491 67.04 17.48
Method in [ho2022classifierfree] (Conditional DDIM) 0.4769 87.93 9.919
Conditional DDIM + $\mathbf{W},\mathbf{C}$ 0.1993 46.89 15.39
@@
To effectively demonstrate the performance and advantages of the proposed I2I algorithm, a comparison in terms of image quality metrics is proposed. Table II displays the values of three image quality/comparison metrics: Frechét Inception Distance (FID), mean Peak Signal-to-Noise Ratio (PSNR), and mean Learned Perceptual Image Patch Similarity (LPIPS). To obtain these values, we use the test set described in Section IV. The FID [Heusel2017] measures the similarity between two data sets of images—usually a data set of generated images and one of reference images—by computing the distance between their feature representations extracted by a pre-trained Inception network. A lower FID indicates better feature alignment of the generated images with the reference data set. The PSNR [Hore2010] measures the reconstruction error between two images in terms of the ratio of the maximum possible pixel value and the mean squared error. A higher PSNR indicates a lower reconstruction error and a better visual quality. It is important to note that PSNR is not very reliable for perceptual quality, as it does not account for the human visual system and the semantic content of the images [Saharia2021]. Nonetheless, in the presented results, we compute the mean PSNR (mPSNR) of all Planet and Synthetic Planet corresponding patches, to further enrich the discussion. The LPIPS [Zhang2018LPIPS] measures the perceptual similarity between a pair of images by computing the distance between their feature representations extracted by a pre-trained deep neural network. A lower LPIPS indicates higher perceptual similarity and better visual quality. LPIPS is aligned with human perception, as it captures the semantic and structural differences between images [Zhang2018]. As with the PSNR, we compute the mean LPIPS (mLPIPS) of all corresponding Planet and Synthetic Planet patches. In a patch-wise comparison, the mLPIPS metric is better suited as a comparison metric than FID since its calculation is done on a patch-by-patch basis instead of focusing on the whole test set at once. Thus, since our tests involve direct comparison between corresponding images, mLPIPS is also employed.
@@
We compare our proposed I2I algorithm with other possible I2I solutions. First, the Sentinel-II test set is directly compared to Planet Dove images without any I2I translation method. The obtained metrics are shown in the first row of Table II. Second, we train two deep regression models [Lathuiliere2020] whose loss objective is to directly match the pixel values of Planet patches when Sentinel-II corresponding patches are input. They share the same neural network backbone as the proposed DDM-based method. In other words, these models are identical to the one used in the proposed method, with the exception that the output linear layer directly predicts the values of pixels of the corresponding Planet patch for a given input. The first deep regression model, whose metrics are displayed in the second row of Table II, is a conventional regression model, i.e., it does not apply any pre or post-processing to its inputs and outputs apart from simple $[-1,1]$ normalization. The second regression model, whose results are displayed in the third row, is identical to the previous model, except that the whitening and coloring algorithms, defined in Algorithm 2 and 4, are applied to its inputs and outputs, respectively. In the table, we refer to this as Regression + $\mathbf{W},\mathbf{C}$. The goal of testing this configuration is to see what would be the influence of these processing methods in a conventional regression model and whether or not it has a significant impact on the metrics. Following, we test the standard classifier-free guided DDIM [ho2022classifierfree], i.e., without the proposed whitening, coloring, and PSNR voting processes, to perform an ablation of these techniques. It is referred to in Table II as Conditional DDIM. Then, we train a similar model, but now adding the whitening and coloring procedures. To measure the impact of the PSNR voting in the inference results, the image quality metrics for this model are separated into two rows, the first being without PSNR voting—denoted as Conditional DDIM + $\mathbf{W},\mathbf{C}$—and the second with it—Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting.
@@
Looking at the numbers of Table II, we observe that the full proposed technique, Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting, reaches, by a good margin, better results in mLPIPS and FID metrics when compared to the other methods. For instance, FID dropped by 52.44 points when compared to Sentinel-II images—an expressive improvement of $53.46\%$. When compared to Regression and Regression + $\mathbf{W}$, the FID has been respectively reduced by 18.48 and by 21.8 points—an improvement of $28.82\%$ and $31.92\%$. Another interesting observation is that the vanilla DDIM could not beat the regression models in any of the metrics. Even worse, it loses to the original images in both mLPIPS and mPSNR. The empirical reason for this is that the vanilla diffusion training seems to get confused by the multiple different overall tonalities of patches containing similar features, which impacts the convergence ability of the model.
@@
This confusion seems to be eliminated when whitening and coloring procedures are used, as evidenced by the performance superiority of both Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. Looking at mLPIPS results, we observe that the regression-based models beat by a small margin the original Sentinel-II, whereas Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting show a much more significant improvement in the same regard. This indicates that, in a patch-by-patch basis, only DDIM + $\mathbf{W},\mathbf{C}$ and DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting models were able to heavily reduce the perceptual distance of patches when compared to their Planet Dove versions. The proposed method with PSNR Voting has reduced mLPIPS by 0.1093 points, meaning that the produced patches are, on average, $36.71\%$ more perceptually similar to their corresponding Planet Dove patches than their Sentinel-II versions. When we compare the last two rows of the table, PSNR voting brings a significant improvement in all metrics: $-5.469\%$ for mLPIPS, $-2.665\%$ for FID, and $+2.144\%$ for mPSNR. We observed, moreover, that PSNR voting is specially useful to discard rare but existent undesired model hallucinations and poor translations, which would negatively impact, for example, change detection results. These improvements, due to the increased number of model evaluations, come at the cost of higher numerical complexity.
@@
While regression models achieve higher mPSNR values, as discussed, PSNR’s preference for blurry regression-produced outputs doesn’t align with human perception. Consequently, learning-based metrics like FID are the standard in image generation comparisons, rendering the PSNR performance superiority of regression-based models irrelevant.

### v2

@@ II-B Inference
The inference procedure, illustrated by Figure 3, aims at maximizing the image generation quality by selecting the best possible initial noise matrix to be fed to the reverse diffusion process. We adapt the classifier-free guided DDIM approach by integrating the same color standardization methods used during training. To further enhance the fidelity of the generated patches to their inputs, we refine the selection process for the initial noise condition input for the reverse diffusion process, ensuring optimal consistency and translation quality. Implementation details for the DDIM inference process are presented in Appendix C-B.
In our experiments, we observed that, for the same conditioning input $\bm{x}$, the quality of DDIM reverse diffusion outputs generated with different initial noise matrices varied significantly. That is, for a combination of independent and identically distributed random initial noise matrices, fixing $\bm{x}$, the DDIM, in some cases, generates samples entirely consistent with $\bm{x}$, whereas sometimes it generates inconsistent samples with undesired hallucinations. Since the DDIM reverse diffusion is a deterministic process, i.e., given fixed $\bm{\varepsilon}_{T}$ and $\bm{x}$, its output remains the same, the cause for varying consistency levels lies only in the different sampled initial noise matrices, which are the source of the randomness of the whole inference process. Thus, we propose a procedure that seeks to select an initial noise matrix such that the DDIM output possesses the highest consistency compared to $\bm{x}$. This procedure consists of repeating the DDIM reverse diffusion process with a reduced number of iterations, using as input $n_{\text{noisy}}$ different initial noise matrices. The resulting outputs are tested for consistency by individually calculating the peak signal-to-noise ratio (PSNR) between them and $\bm{x}_{\text{local}}$. The one that was responsible for the maximum PSNR is used as the definitive starting noise matrix for the inference process, which takes place with a higher number of iterations to ensure maximal generation quality.
@@
III Dataset
@@
For the proposed experiments, paired images from satellites Sentinel-II and Planet Dove were gathered. The former is a widely known open-access optical satellite with a spatial resolution of 10 meters, whereas the latter is a class of commercial satellites aimed at producing high-quality images at a spatial resolution of 3 meters. A dataset containing 77 pairs of surface reflectance rasters has been assembled for training. These rasters are parts of or a combination of acquisitions that were captured at similar time periods. They are properly geolocated, georeferenced, and checked for image quality issues. In total, a vast area of 7894 $7894\text{ km}^{2}$ is covered by the training set, which is comparable to the area of Puerto Rico (9104 $9104\text{ km}^{2}$). As the test set for the image-to-image translation algorithm, 8 pairs of images were selected, which, in total, cover an area of 707 $707\text{ km}^{2}$. With these, image quality metrics are extracted for the proposed method and compared to other possible approaches for the I2I translation task.
@@
For evaluation of the performance of our method in a change detection task, we use pre and post-images from the port region in Beirut, Lebanon. On August 4th, 2020, an explosion dramatically affected the port and surrounding areas. As they have recovered over time, many changes took place, e.g., new and repaired constructions and structures. Thus, for the change detection test, we use as pre-event images a pair of images acquired hours after the explosion, between 04 and 20 of August 2020, when the effects of the event are still evident. As post images, acquisitions from July 2023 of the exact same region are used. Figure 1 shows the Beirut pre and post-event images from Sentinel-II and Planet Dove. We also perform change detection tests in a region in Austin, USA, where many urban changes can be observed between the selected pre- and post-event images. They were captured in July 2021 and 2023, respectively.
@@
(b) Planet Dove (Post)
(c) Proposed: DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting
(d) DDIM + $\mathbf{W},\mathbf{C}$
(e) DDIM
(f) Regression
(g) Regression + $\mathbf{W},\mathbf{C}$
(h) Pix2Pix [Pix2Pix2017]
(i) ShuffleMixer [sun2022shufflemixer]
(j) SRDenseNet [zhang2018residual]
(k) SwinIR [liang2021swinir]
(l) SR3 [Saharia2021SR3]
Figure 4: Comparison among images generated by the tested I2I methods (c)-(l) and the original Sentinel-II (a) and Planet Dove (b) images. The image generated with the proposed model (c) displays higher resolution features not observable among the others, while maintaining patch-wise feature consistency and avoiding blurriness.
IV Experiments
IV-A Image Quality Metrics and Ablation Experiments
TABLE I: Image Quality Metrics Comparison with the State-of-the-Art Using $N_{\text{final}}=64$, $N_{\text{pre}}=8$, $n_{\text{noisy}}=8$
@@ IV-A Image Quality Metrics and Ablation Experiments
Sentinel-II (Input Images) 0.2977 98.08 12.76
Regression 1 0.2797 64.12 16.72
Regression + $\mathbf{W},\mathbf{C}$ 1 0.2491 67.04 17.48
Conditional DDIM 1 0.4769 87.93 9.919
Conditional DDIM + $\mathbf{W},\mathbf{C}$ 1 0.1993 46.89 15.39
@@
Other State of the Art Models
Pix2Pix [Pix2Pix2017] 0.3211 105.0 14.87
ShuffleMixer [sun2022shufflemixer] 0.2249 73.82 16.06
SRDenseNet [zhang2018residual] 0.2296 57.70 17.12
SwinIR [liang2021swinir] 0.2325 56.36 16.68
SR3 [Saharia2021SR3] 0.6235 152.6 9.906
1 Identical neural network architecture as the proposed method.
TABLE II: Comparison Metrics for Different Inference Hyperparameters
[…]
16 2 2 20 0.2229 52.46 14.60
32 4 4 48 0.1967 45.17 15.29
48 6 6 84 0.2013 46.49 15.15
64 8 8 128 0.1884 45.64 15.72
@@
To effectively demonstrate the performance and advantages of the proposed I2I algorithm, a comparison in terms of image quality metrics is proposed. Table I displays the values of three image quality/comparison metrics: Frechét Inception Distance (FID), mean Peak Signal-to-Noise Ratio (PSNR), and mean Learned Perceptual Image Patch Similarity (LPIPS). To obtain these values, we use the test set described in Section III. The FID [Heusel2017] measures the similarity between two data sets of images—usually a data set of generated images and one of reference images—by computing the distance between their feature representations extracted by a pre-trained Inception network. A lower FID indicates better feature alignment of the generated images with the reference data set. The PSNR [Hore2010] measures the reconstruction error between two images in terms of the ratio of the maximum possible pixel value and the mean squared error. A higher PSNR indicates a lower pixel-wise reconstruction error. It is important to note that PSNR is not a reliable metric for perceptual quality, as it does not account for the human visual system and the semantic content of the images [Saharia2021]. Nonetheless, in the presented results, we compute the mean PSNR (mPSNR) of all Planet and Synthetic Planet corresponding patches, to further enrich the discussion. The LPIPS [Zhang2018LPIPS] measures the perceptual similarity between a pair of images by computing the distance between their feature representations extracted by a pre-trained deep neural network. A lower LPIPS indicates higher perceptual similarity and better visual quality. LPIPS is aligned with human perception, as it captures the semantic and structural differences between images [Zhang2018LPIPS]. As with the PSNR, we compute the mean LPIPS (mLPIPS) of all corresponding Planet and Synthetic Planet patches. In a patch-wise comparison, the mLPIPS metric is better suited as a comparison metric than FID since its calculation is done on a patch-by-patch basis instead of focusing on the whole test set at once. Thus, since our tests involve direct comparison between corresponding images, mLPIPS is also employed.
@@
We compare our proposed I2I algorithm with other possible I2I solutions. First, the Sentinel-II test set is directly compared to Planet Dove images without any I2I translation method. The obtained metrics are shown in the first row of Table I. Second, we train two deep regression models [Lathuiliere2020] whose loss objective is to directly match the pixel values of Planet patches when Sentinel-II corresponding patches are input. They share the same neural network backbone as the proposed DDM-based method. In other words, these models are identical to the one used in the proposed method, with the exception that the output linear layer directly predicts the values of pixels of the corresponding Planet patch for a given input. The first deep regression model, whose metrics are displayed in the second row of Table I, is a conventional regression model, i.e., it does not apply any pre or post-processing to its inputs and outputs apart from simple $[-1,1]$ normalization. The second regression model, whose results are displayed in the third row, is identical to the previous model, except that the whitening and coloring algorithms, defined in Algorithm 2 and 4, are applied to its inputs and outputs, respectively. In the table, we refer to this as Regression + $\mathbf{W},\mathbf{C}$. The goal of testing this configuration is to see what would be the influence of these processing methods in a conventional regression model and whether or not it has a significant impact on the metrics. Following, we test the standard classifier-free guided DDIM [ho2022classifierfree], i.e., without the proposed whitening, coloring, and PSNR voting processes, using the exact same model architecture as the proposed method, to perform an ablation of these techniques. It is referred in Table I as Conditional DDIM. Then, we train an identical model, now adding the whitening and coloring procedures. To measure the impact of the PSNR voting in the inference results, the image quality metrics for this model are separated into two rows, the first being without PSNR voting—denoted as Conditional DDIM + $\mathbf{W},\mathbf{C}$—and the second with it—Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. All DDIM results were obtained using $N_{\text{final}}=64$. The proposed method, which includes PSNR voting, used $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$.
We also compare the performance of our method with five other state-of-the-art approaches for super resolution and image-to-image translation tasks. The first method is the well-established Pix2Pix [Pix2Pix2017], which has been proposed for multiple generative tasks, including image-to-image translation. It is an adversarial framework that uses both regression and discriminative losses to guide training of a generator UNet-like network. Following, we compare our method against the SwinIR [liang2021swinir] super resolution model, which leverages the Swin Transformer architecture for enhanced image processing capabilities. The ShuffleMixer [sun2022shufflemixer] architecture, known for its efficient mixing of features through channel shuffling and spatial mixing, is another model we compare against. The SRDenseNet [zhang2018residual] is also included in our comparison—a model that utilizes dense connections to improve the flow of information and gradients throughout the network, which is particularly beneficial for the task of super resolution. Finally, we present results for the SR3 super resolution diffusion-based framework [Saharia2021SR3], using DDIM sampling with 64 steps. The UNet structure of SR3 is analogous to the one used in the article that introduced the DDPM procedure [Ho2020]. Thus, comparing it to our method provides an insight on the advantages of the neural network structure chosen for our experiments.
IV-A1 Comparison Metrics
@@
Looking at the numbers of Table I, we observe that the full proposed technique, Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting, reaches, by a good margin, better results in mLPIPS and FID metrics when compared to all other methods and to the Sentinel-II baseline. FID dropped by 52.44 points when compared to Sentinel-II images—an expressive improvement of $53.46\%$. When compared to Regression and Regression + $\mathbf{W}$, the FID has been respectively reduced by 18.48 and by 21.8 points—an improvement of $28.82\%$ and $31.92\%$. Among the state-of-the-art methods, only SwinIR and SRDenseNet reached comparable mLIPS and FID values to the ones from our approach.
@@
Interestingly, the vanilla DDIM-based models (Conditional DDIM and SR3) lose to all methods in any of the metrics. Even worse, they lose to the original images in both mLPIPS and mPSNR. The empirical reason for this is that the vanilla diffusion training seems to get confused by the multiple different overall tonalities of patches containing similar features, which impacts the convergence ability of the model. Deck and Bischoff [deck2023easing] noticed that diffusion models tend to present such color shifting problems, which tends to get worse for larger images. This issue seems to be eliminated when whitening and coloring procedures are used, as evidenced by the performance superiority of both Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting. Looking at mLPIPS results, we observe that the regression-based models beat by a small margin the original Sentinel-II images, whereas Conditional DDIM + $\mathbf{W},\mathbf{C}$ and Conditional DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting show a much more significant improvement in the same regard. This indicates that, in a patch-by-patch basis, only DDIM + $\mathbf{W},\mathbf{C}$ and DDIM + $\mathbf{W},\mathbf{C}$ + PSNR Voting models were able to heavily reduce the perceptual distance of patches when compared to their Planet Dove versions. The proposed method with PSNR Voting has reduced mLPIPS by 0.1093 points, meaning that the produced patches are, on average, $36.71\%$ more perceptually similar to their corresponding Planet Dove patches than their Sentinel-II versions. When we compare the last two rows of the table, PSNR voting brings a significant improvement in all metrics: $-5.469\%$ for mLPIPS, $-2.665\%$ for FID, and $+2.144\%$ for mPSNR. We observed, moreover, that PSNR voting is specially useful to discard rare but existent undesired model hallucinations and poor translations, which would negatively impact, for example, change detection results. These improvements, due to the increased number of model evaluations, come at the cost of higher numerical complexity.
@@
While regression models achieve higher mPSNR values, as discussed, PSNR’s preference for blurry regression-produced outputs does not align well with human perception. The same can be said about the SwinIR, ShuffleMixer and SRDenseNet methods, as they are also trained using regression objectives. Consequently, learning-based metrics, like the FID and the LPIPS, lately have been the standard in image generation comparisons, rendering the PSNR performance superiority of regression models irrelevant.
As discussed, the very poor performance shown by the SR3 model is in part due to what plagued the performance of the classically trained DDIM model. However, even compared to it, the SR3 model presented far worse numbers. This fact shows how the model architecture chosen for our method has been able to better extract the complex features of the conditioning input and generate better translations.
IV-A2 Inference Hyperparameters
To verify the impact of the choice of inference hyperparameters linked to the proposed method, we present a performance comparison for different values of $N_{\text{final}}$, $N_{\text{pre}}$ and $n_{\text{noisy}}$ in Table II. In terms of mLPIPS, $N_{\text{final}}=64$, $N_{\text{pre}}=8$ and $n_{\text{noisy}}=8$ reached the best result of $0.1884$. We use this configuration in our change detection experiments, as a lower perceptual distance between identical patches is crucial for a high change detection performance. Interestingly, however, $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ reached a lower FID, which suggests that further increasing the number of iterations does not bring any value with regards to image quality. Moreover, unintuitively, $N_{\text{final}}=48$, $N_{\text{pre}}=6$ and $n_{\text{noisy}}=6$ showed worse performance metrics compared to the former configuration. This may be due to the fact that the number of training diffusion steps was set to 1024, and 48 is not a divisor of such number. This implicates that the particular timestep values input to the model during the DDIM sampling were not seen during training, leading to such worse performance values. Finally, for $N_{\text{final}}=16$, $N_{\text{pre}}=2$ and $n_{\text{noisy}}=2$, the performance metrics were clearly degraded, which suggests that $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ would be a good balance between numerical complexity and performance. In addition, the usefulness of PSNR voting is confirmed when this configuration is compared to the results of the model using 64 DDIM iterations without PSNR voting, displayed in Table I. That is, even with a lower total number of iterations (48 vs 64), the configuration $N_{\text{final}}=32$, $N_{\text{pre}}=4$ and $n_{\text{noisy}}=4$ reached better mLPIPS and FID values compared to the DDIM without PSNR voting.